# Who Wrote this Code? Watermarking for Code Generation

### v3

@@ Abstract
Since the remarkable generation performance of large language models raised ethical and legal concerns, approaches to detect machine-generated text by embedding watermarks are being developed. However, we discover that the existing works fail to function appropriately in code generation tasks due to the task’s nature of having low entropy. Extending a logit-modifying watermark method, we propose Selective WatErmarking via Entropy Thresholding (SWEET), which enhances detection ability and mitigates code quality degeneration by removing low-entropy segments at generating and detecting watermarks. Our experiments show that SWEET significantly improves code quality preservation while outperforming all baselines, including post-hoc detection methods, in detecting machine-generated code text. Our code is available in the supplementary materials.
@@
[…]
footnotetext: $\star$ Authors equally contributed.footnotetext: $\sharp$ Work done during Ilgee’s internship at SNU VL Lab.footnotetext: $\dagger$ Corresponding authors.
@@ 3.1 Motivation
Low Entropy Sequences Avoid Being Watermarked. Another critical issue is when watermark strength is too weak to embed watermarks into a text with low entropy. If a red list token has a too high logit value to be inevitably generated, it hinders watermark detection. For example, in Figure 2 (a), tokens with white backgrounds representing low entropy have few green tokens. This becomes much more fatal in code generation tasks where outcomes are relatively shorter than the plain text, such as asking only a code block of a function [^6] The average token length of human-written solution codes in HumanEval, MBPP, and DS-1000 datasets is only 57.. The WLLM detection method is based on a statistical test, which involves counting the number of green list tokens in the entire length. However detecting watermarks based on a statistical test deteriorates if the length is short. [^7] We measured detectability according to the length of generated texts and observed that WLLM performs relatively poorly while SWEET is robust in detecting watermarks within short texts. For more details, please refer to Appendix E.
@@ 3.3 Effect of Entropy Thresholding
WLLM ($\Delta$pass@1 $\sim-10\%$ ${}^{\star}$ 29.6 0.822 0.402 < 0.05 34.5 0.718 0.178 < 0.05 23.9 0.627 0.152 < 0.05
SWEET ($\Delta$pass@1 $\sim-10\%$ ${}^{\star}$ 32.6 0.943 0.835 < 0.05 33.8 0.873 0.590 < 0.05 23.7 0.815 0.384 < 0.05
WLLM (AUROC0.9absent0.9$\geq$ 0.9≥ ${}^{\dagger}$ 25.3 0.904 0.652 < 0.05 24.2 0.930 0.718 < 0.05 8.6 0.944 0.793 < 0.05
SWEET (AUROC0.9absent0.9$\geq$ 0.9≥ ${}^{\dagger}$ 32.6 0.943 0.835 < 0.05 33.2 0.906 0.548 < 0.05 18.8 0.924 0.649 < 0.05
Table 1: Main results of code generation performance and detection ability. Since calibration on watermarking strength leads to trade-offs between code generation quality and detection ability, we present two results for WLLM and SWEET. ${}^{\star}$ for the best detection score (i.e., AUROC and TPR) while allowing a code generation quality decrease of $\sim$10compared to Non-watermarked, and ${}^{\dagger}$ for the best code generation quality (pass@1) among AUROC $\geq$ 0.9. The selected points are shown in Figure 3. We add EXP-edit and a Non-watermarked baseline with a high entropy setting (i.e., temperature=1.0 and top-p=1.0) since EXP-edit hardly detects watermarking in low entropy environments.
Figure 3: The tradeoff between AUROC and pass@1 of detecting real and generated samples of HumanEval, MBPP, and DS-1000 datasets. The pink line represents a Pareto frontier of SWEET, while the blue line represents that of WLLM. SWEET shows consistent dominance. The red/orange line and circles are the points used in Table 1. The entropy threshold value used for SWEET is 1.2 here, and Pareto frontier figures for all threshold values are in Figure 6.
@@ 4.1 Tasks and Metrics
We select three python code generation tasks, HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and DS-1000 Lai et al. (2023), as our testbeds. These tasks contain python programming problems, test cases, and human-written canonical answers. Language model is prompted with programming problems and expected to generate the correct code that can pass the test cases.
@@ 5.1 Main Results
Detection Performance. Table 1 shows that overall, our SWEET method outperforms all baselines in detecting machine-generated code with a price of 10% degradation of code functionality. Both in the MBPP and DS-1000 datasets, SWEET achieves AUROC of 0.873 and 0.815, respectively, whereas none of the baselines exceeded 0.8. SWEET even achieves an AUROC of 0.943 in HumanEval with a 2.4% degradation of code functionality. However, when only near 10% degradation of code functionality is allowed, WLLM shows lower detection performance than our method. In the case of the distortion-free watermarking method, due to the lower entropy of the code generation task, EXP-edit fails to achieve an AUROC score exceeding 0.6 in all cases, and even EXP-edit with high entropy setting could not outperform our methods with regard of the detection performance. While all post-hoc detection baselines preserve code functionality as they do not modify generated code, none of them achieve an AUROC score above 0.6.999We defer a more in-depth discussion about the breakdown of Post-hoc methods to Appendix J.
@@ 6.1 Impact of Entropy Thresholds
Figure 4 presents how code generation performance and detecting ability trade-off when calibrating the entropy threshold in our method. WLLM is when the entropy threshold is not applied (i.e., entropy threshold=0). As the entropy threshold increases, the ratio of watermarked tokens decreases, so the code generation performance converges to a non-watermarked base model. This indicates that our method always lies between the WLLM and a non-watermarked base model in terms of code generation performance. On the other hand, the detection ability, as the entropy threshold increases, reaches a local maximum but eventually declines. While our method with a moderate threshold effectively restricts generating the red list tokens compared to the WLLM, detection ability eventually decreases if the threshold is so high that few tokens are watermarked.
@@ 6.2 Detection Ability without Prompts
As entropy information is required in the detection phase, approximating entropy values for each generation time step $t$ is essential in our method. In the main experiments, we prepend the prompt used in the generation phase (e.g., the question of Fig. 2) before the target code to reproduce the same entropy. However, we hardly know the prompt used for a given target code in the real world. Thus, instead of using the gold prompt, we attach a common and general prompt for code generation to approximate the entropy information, such as "def solution(*args): """Generate a solution"""". We use five general prompts (see Appendix G), and their z-scores are averaged for use in detection.
@@ 6.3 Use of Surrogate Model
In the results of Figure 9, the detection performance declines are insignificant, and our approach utilizing the surrogate model continues to surpass the baseline. Such performance preservation may be due to that LLaMA2 7B and 13B are trained on the identical training corpus (Touvron et al., 2023). Further analysis for computational cost can be found in Appendix H.
6.4 Robustness to Variable Renaming
Even with the text watermarked, a malicious user might attempt to remove watermarks in the text by paraphrasing (Krishna et al., 2023; Sadasivan et al., 2023). Paraphrasing the code text is more restrictive than dealing with plain text because it must avoid triggering any code malfunctions. We assess the robustness of watermarking methods against paraphrasing by employing a straightforward approach - changing the names of variables. Specifically, we select variables in the watermarked code and rename them with randomly generated strings of varying lengths, ranging from 2 to 5 characters.
Figure 10 presents the results of the detection performance on the renamed variables in the code. All watermarking methods show the decline of AUROC scores when more variables are renamed, while our approaches continue to show better performances than baselines. However, our approaches also show that the AUROC scores drop to about 0.8 when all variables are renamed. We found that this is because variable names comprise a large proportion of high entropy tokens in the code text (See Appendix I for details).
@@ 7 Conclusion
We identify limitations of this work and suggest ways to mitigate them. We want to note that these limitations are not weaknesses of our work as they are shared by the status quo of this field.
The first one is about the robustness issue. As users can tailor LLM’s code to their specific needs, it is crucial to be robust against any kind of paraphrasing attacks. Though such robustness is not the focus of this work, we acknowledge the importance of robustness and leave it for future work.
The necessity to calibrate an additional hyperparameter, the entropy threshold, for our method could pose another limitation. While we demonstrate that our method with a moderate threshold value outperforms the baselines (see Sec 6.1), the imperative manual procedure of selecting the entropy threshold demands additional computational expenses.
Furthermore, during detection, we also need the source Code LLM, hence this method works only in a completely white-box setting. Although it has been shown that employing even a smaller surrogate LM can still maintain the detection performances to some degree (see Sec 6.3), this can be a computational burden for some users who want to apply our work.
@@ Appendix C Proof of Theorem 1
WLLM ($\Delta$pass@1 $\sim-10\%$ ${}^{\star}$ 15.4 0.777 0.402 <0.05
SWEET ($\Delta$pass@1 $\sim-10\%$ ${}^{\star}$ 15.5 0.921 0.616 <0.05
WLLM (AUROC0.9absent0.9$\geq$ 0.9≥ ${}^{\dagger}$ 9.2 0.908 0.720 <0.05
SWEET (AUROC0.9absent0.9$\geq$ 0.9≥ ${}^{\dagger}$ 15.5 0.921 0.616 <0.05
@@ Appendix D Implementation Details
We have used three datasets for our testbeds: HumanEval, MBPP, and DS-1000. They have 164, 500, and 1000 Python code problems, respectively. For our base models, StarCoder and LLaMA2, we use top-$p$ (Holtzman et al., 2020) sampling with $p=0.95$ for both models, and temperature 0.2 and 0.1, respectively. When generating output for each code problems, we use zero-shot setting in HumanEval and DS-1000 but 3-shot in MBPP. Prompts used in MBPP are similar to the prompt in Austin et al. (2021). For calculating pass@1 scores, we set $n=40$ for HumanEval and DS-1000, and $n=20$ for MBPP.
@@ D.3 EXP-edit
Appendix E Detectability with Varying Code Lengths

### v4

@@ Abstract
Since the remarkable generation performance of large language models raised ethical and legal concerns, approaches to detect machine-generated text by embedding watermarks are being developed. However, we discover that the existing works fail to function appropriately in code generation tasks due to the task’s nature of having low entropy. Extending a logit-modifying watermark method, we propose Selective WatErmarking via Entropy Thresholding (SWEET), which enhances detection ability and mitigates code quality degeneration by removing low-entropy segments at generating and detecting watermarks. Our experiments show that SWEET significantly improves code quality preservation while outperforming all baselines, including post-hoc detection methods, in detecting machine-generated code text. Our code is available in https://github.com/hongcheki/sweet-watermark.
@@
 Taehyun Lee⋆,1  Seokhee Hong⋆,1,3  Jaewoo Ahn1  Ilgee Hong1,4,♯ Hwaran Lee2Sangdoo Yun1,2Jamin Shin†,2,♮Gunhee Kim†,1 1Seoul National University 2NAVER AI Lab 3LG AI Research  4Georgia Institute of Technology {taehyun.lee, seokhee.hong, jaewoo.ahn}@vision.snu.ac.kr ihong39@gatech.edu, {hwaran.lee, sangdoo.yun}@navercorp.com jamin.shin@outlook.com, gunhee@snu.ac.kr
footnotetext: $\star$ Authors equally contributed.footnotetext: $\sharp$ Work done during Ilgee’s internship at SNU VL Lab.footnotetext: $\natural$ Jamin Shin is currently affiliated with Trillion Labs, but the work was done while he was at NAVER. footnotetext: $\dagger$ Corresponding authors.
@@ 3.1 Motivation
Low Entropy Sequences Avoid Being Watermarked. Another critical issue is when watermark strength is too weak to embed watermarks into a text with low entropy. If a red list token has a too high logit value to be inevitably generated, it hinders watermark detection. For example, in Figure 2 (a), tokens with white backgrounds representing low entropy have few green tokens. This becomes much more fatal in code generation tasks where outcomes are relatively shorter than the plain text, such as asking only a code block of a function [^6] The average token length of human-written solution codes in HumanEval, MBPP, and DS-1000 datasets is only 57.. The WLLM detection method is based on a statistical test, which involves counting the number of green list tokens in the entire length. However detecting watermarks based on a statistical test deteriorates if the length is short. [^7] We measured detectability according to the length of generated texts and observed that WLLM performs relatively poorly while SWEET is robust in detecting watermarks within short texts. For more details, please refer to Appendix G.
@@ 3.3 Effect of Entropy Thresholding
WLLM ($\Delta$pass@1 $\sim-10\%$) 29.6 0.822 0.402 < 0.05 34.5 0.718 0.178 < 0.05 23.9 0.627 0.152 < 0.05
SWEET ($\Delta$pass@1 $\sim-10\%$) 32.6 0.943 0.835 < 0.05 33.8 0.873 0.590 < 0.05 23.7 0.815 0.384 < 0.05
WLLM (AUROC0.9absent0.9$\geq$ 0.9≥ 0.9) 25.3 0.904 0.652 < 0.05 24.2 0.930 0.718 < 0.05 8.6 0.944 0.793 < 0.05
SWEET (AUROC0.9absent0.9$\geq$ 0.9≥ 0.9) 32.6 0.943 0.835 < 0.05 33.2 0.906 0.548 < 0.05 18.8 0.924 0.649 < 0.05
Table 1: Main results of code generation performance and detection ability. Since calibration on watermarking strength leads to trade-offs between code generation quality and detection ability, we present two results for WLLM and SWEET. for the best detection score (i.e., AUROC and TPR) while allowing a code generation quality decrease of $\sim$10compared to Non-watermarked, and for the best code generation quality (pass@1) among AUROC $\geq$ 0.9. The selected points are shown in Figure 3. We add EXP-edit and a Non-watermarked baseline with a high entropy setting (i.e., temperature=1.0 and top-p=1.0).
@@ 4.1 Tasks and Metrics
We select three Python code generation tasks, HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and DS-1000 Lai et al. (2023), as our main testbeds. These tasks contain Python programming problems, test cases, and human-written canonical answers. Language models are prompted with programming problems and expected to generate the correct code that can pass the test cases. To evaluate our approach’s performance in more diverse software development contexts, such as other languages or other code generation scopes, we also include two more datasets: HumanEvalPack (Muennighoff et al., 2024) and ClassEval (Du et al., 2023). Please refer to Appendix E for implementation details of these benchmarks.
@@ 5 Results
Method HumanEvalPack - C++ HumanEvalPack - Java ClassEval
pass@1 AUROC TPR FPR pass@1 AUROC TPR FPR pass@5 AUROC TPR FPR
Non-watermarked 29.4 - - - 31.5 - - - 14.0 - - -
Non-watermarked (w/ high entropy) 18.2 - - - 13.9 - - - 19.0 - - -
Post-hoc
log p(x) 29.4 0.656 0.160 < 0.05 31.5 0.635 0.127 < 0.05 14.0 0.847 0.320 < 0.05
LogRank 0.658 0.187 < 0.05 0.654 0.240 < 0.05 0.821 0.260 < 0.05
DetectGPT (T5-3B) 0.646 0.079 < 0.05 0.699 0.273 < 0.05 0.610 0.140 < 0.05
DetectGPT 0.525 0.079 < 0.05 0.650 0.116 < 0.05 0.749 0.210 < 0.05
GPTZero 0.486 0.073 < 0.05 0.529 0.000 < 0.05 0.885 0.800 < 0.05
OpenAI Classifier 0.631 0.120 < 0.05 0.545 0.087 < 0.05 0.503 0.010 < 0.05
Watermarking
EXP-edit 28.3 0.605 0.091 < 0.05 32.1 0.486 0.024 < 0.05 21.0 0.497 0.020 < 0.05
EXP-edit (w/ high entropy) 16.7 0.749 0.402 < 0.05 14.3 0.828 0.512 < 0.05 21.0 0.513 0.040 < 0.05
WLLM ($\Delta$pass@1 $\sim-10\%$) 25.9 0.887 0.604 < 0.05 25.5 0.833 0.518 < 0.05 12.0 0.939 0.840 < 0.05
SWEET ($\Delta$pass@1 $\sim-10\%$) 26.2 0.943 0.817 < 0.05 27.6 0.862 0.457 < 0.05 13.0 0.980 0.920 < 0.05
WLLM (AUROC0.9absent0.9$\geq$ 0.9≥ 0.9) 25.9 0.887 0.604 < 0.05 9.5 0.947 0.872 < 0.05 12.0 0.939 0.840 < 0.05
SWEET (AUROC0.9absent0.9$\geq$ 0.9≥ 0.9) 29.0 0.904 0.707 < 0.05 22.6 0.969 0.878 < 0.05 13.0 0.980 0.920 < 0.05
Table 2: Main results of code generation performance and detection ability on HumanEvalPack (Muennighoff et al., 2024) and ClassEval (Du et al., 2023). Since calibration on watermarking strength leads to trade-offs between code generation quality and detection ability, we present two results for WLLM and SWEET. for the best detection score (i.e., AUROC and TPR) while allowing a code generation quality decrease of $\sim$10compared to Non-watermarked, and for the best code generation quality (pass@1) among AUROC $\geq$ 0.9. We add EXP-edit and a Non-watermarked baseline with a high entropy setting (i.e., temperature=1.0 and top-p=1.0).
Figure 3: The tradeoff between AUROC and pass@1 of detecting real and generated samples of HumanEval, MBPP, and DS-1000 datasets. The pink line represents a Pareto frontier of SWEET, while the blue line represents that of WLLM. SWEET shows consistent dominance. The red/orange line and circles are the points used in Table 1. The entropy threshold for SWEET is 1.2 here, and Pareto frontier figures for all threshold values are in Figure 6.
@@ 5.1 Main Results
Detection Performance. Table 1 shows that overall, our SWEET method outperforms all baselines in detecting machine-generated code with a price of 10% degradation of code functionality. Both in the MBPP and DS-1000 datasets, SWEET achieves AUROC of 0.873 and 0.815, respectively, whereas none of the baselines exceeded 0.8. SWEET even achieves an AUROC of 0.943 in HumanEval with a 2.4% degradation of code functionality. However, when only near 10% degradation of code functionality is allowed, WLLM shows lower detection performance than our method. In the case of the distortion-free watermarking method, due to the lower entropy of the code generation task, EXP-edit fails to achieve an AUROC score exceeding 0.6 in all cases, and even EXP-edit with high entropy setting could not outperform our methods with regard of the detection performance. While all post-hoc detection baselines preserve code functionality as they do not modify generated code, none of them achieve an AUROC score above 0.6.999We defer a more in-depth discussion about the breakdown of Post-hoc methods to Appendix K.
@@
C++/Java/Class-level Code Generation. Table 2 presents results on other programming languages (C++ and Java) and another code generation scope (i.e., class-level). While preserving code functionality much more than WLLM, SWEET shows the highest detection performance except in the Java environment, where the TPR score of WLLM is higher than that of SWEET. The results demonstrate that the efficacy of our methodology is not limited to certain types of programming languages or software development environments. For more analysis of the results, please refer to Appendix E.
@@ 6.1 Impact of Entropy Thresholds
Figure 4 presents how code generation performance and detecting ability trade-off when calibrating the entropy threshold in our method. WLLM is when the entropy threshold is not applied (i.e., entropy threshold=0). As the entropy threshold increases, the ratio of watermarked tokens decreases, so the code generation performance converges to a non-watermarked base model. This indicates that our method always lies between the WLLM and a non-watermarked base model in terms of code generation performance. On the other hand, the detection ability, as the entropy threshold increases, reaches a local maximum but eventually declines. While our method with a moderate threshold effectively restricts generating the red list tokens compared to the WLLM, detection ability eventually decreases if the threshold is so high that few tokens are watermarked. We further investigate how to effectively calibrate the entropy threshold value in Appendix H.
@@ 6.2 Detection Ability without Prompts
As entropy information is required in the detection phase, approximating entropy values for each generation time step $t$ is essential in our method. In the main experiments, we prepend the prompt used in the generation phase (e.g., the question of Fig. 2) before the target code to reproduce the same entropy. However, we hardly know the prompt used for a given target code in the real world. Thus, instead of using the gold prompt, we attach a common and general prompt for code generation to approximate the entropy information. We use five general prompts as below, and their z-scores are averaged for use in detection.
def solution(*args):
Generate a solution
<filename>solutions/solution_1.py
# Here is the correct implementation of the code exercise
def solution(*args):
def function(*args, **kargs):
Generate a code given the condition
from typing import List
def my_solution(*args, **kargs):
Generate a solution
def foo(*args):
Solution that solves a problem
@@ 6.3 Use of Surrogate Model
In the results of Figure 10, the detection performance declines are insignificant, and our approach utilizing the surrogate model continues to surpass the baseline. Such performance preservation may be due to that LLaMA2 7B and 13B are trained on the identical training corpus (Touvron et al., 2023). Further analysis for computational cost can be found in Appendix I.
6.4 Robustness to Paraphrasing Attacks
Figure 5: Watermark detection performance on renamed variables in the code. We set $\gamma=0.25$ and $\delta=3.0$ for WLLM and SWEET. For EXP-edit, we search the hyperparameter for the block size in [20,30,40] with a high entropy setting.
Even with the text watermarked, a malicious user might attempt to remove watermarks in the text by paraphrasing (Krishna et al., 2023; Sadasivan et al., 2023). Paraphrasing the code text is more restrictive than dealing with plain text because it must avoid triggering any code malfunctions. We assess the robustness of watermarking methods against paraphrasing by employing two types of attacks - changing the names of variables and utilizing a commercial code refactoring service. [^10] https://codepal.ai/code-refactor Specifically, for each watermark method, we choose 273 source codes from the MBPP task, for which all three methods succeed in generating with no syntax error. In the code renaming attack, we select variables in the watermarked code and rename them with randomly generated strings of varying lengths, ranging from 2 to 5 characters. We use five random seeds for renaming.
Figure 5 presents the results of the detection performance on the paraphrased code. All watermarking methods show the decline of AUROC scores when the extent of paraphrasing increases, while our approaches continue to show better performances than baselines. However, our approaches also show that the AUROC scores drop to about 0.8 when all variables are renamed. We found that this is because variable names comprise a large proportion of high entropy tokens in the code text (See Appendix J for details).
@@ 7 Conclusion
We identify the limitations of this work and suggest ways to mitigate them. First, two issues are shared by the status quo of this field as follows. (1) Robustness against paraphrasing attacks: As users can tailor LLM’s code to their specific needs, it is crucial to be robust against paraphrasing attacks. We addressed this issue in Section 6.4 and left further robustness enhancement for future work. (2) Possibilities of watermark forgery: An attacker may pry out the watermarking rules, and $\mathcal{O}(|\mathcal{V}|)^{2}$ runs in the brute-force mechanism enable it. Against the attack, one can apply techniques enhancing the watermarking model’s security, such as dividing the green/red list depending on prior $h>1$ tokens, as mentioned in the WLLM paper, or applying methods like SelfHash Kirchenbauer et al. (2023b).
For our work, two additional issues exist as follows. (1) Entropy threshold calibration: We demonstrate that our method outperforms the baselines in the broad entropy threshold range (see Sec 6.1) and investigate how to calibrate the entropy threshold effectively (see Appendix H). However, we still need entropy threshold tuning to obtain the best performance, which costs a computation. (2) Need for the source LLM during detection: SWEET works in a white-box setting. Although it has been shown that employing even a smaller surrogate LM can still maintain the detection performances to some degree (see Sec 6.3), this can be a computational burden for some users who want to apply our work.
@@
Acknowledgements
We would like to thank the anonymous reviewers for their valuable feedback. This work was supported by SNU-NAVER Hyperscale AI Center, Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2019-II191082, SW StarLab), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2023R1A2C2005573), and the Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education(RS-2023-00274280).
@@ Appendix C Proof of Theorem 1
WLLM ($\Delta$pass@1 $\sim-10\%$) 15.4 0.777 0.402 <0.05
SWEET ($\Delta$pass@1 $\sim-10\%$) 15.5 0.921 0.616 <0.05
WLLM (AUROC0.9absent0.9$\geq$ 0.9≥ 0.9) 9.2 0.908 0.720 <0.05
SWEET (AUROC0.9absent0.9$\geq$ 0.9≥ 0.9) 15.5 0.921 0.616 <0.05
@@ Appendix D Implementation Details
We have used three datasets for our testbeds: HumanEval, MBPP, and DS-1000. They have 164, 500, and 1000 Python code problems, respectively. For our base models, StarCoder and LLaMA2, we use top-$p$ (Holtzman et al., 2020) sampling with $p=0.95$ for both models, and temperature 0.2 and 0.1, respectively. When generating output for each code problems, we use zero-shot setting in HumanEval and DS-1000 but 3-shot in MBPP. Prompts used in MBPP are similar to the prompt in Austin et al. (2021). For calculating pass@1 scores, we set $n=40$ for HumanEval and DS-1000, and $n=20$ for MBPP. We use a single NVIDIA RTX A6000 GPU to generate or detect each code completion with StarCoder or LLaMA2. It takes less than two GPU hours for generation and less than 1 GPU hour for detection.
@@ D.3 EXP-edit
Appendix E Experimental Details and Results on HumanEvalPack and ClassEval
We choose two additional benchmarks to present how well our approach functions in broader software development contexts.
E.1 HumanEvalPack
HumanEvalPack (Muennighoff et al., 2024) is an extension of HumanEval to cover 6 languages (Python, C++, Java, Javascript, Go, and Rust), and we choose C++ and Java as our testbeds. Basically, all hyperparameter settings are equal to Python benchmarks except $n=5$. For WLLM and SWEET, we narrow the search space for $\delta$ into [1.0,2.0,3.0,4.0]. For EXP-edit, we fix the length of the key sequence to 100.
E.2 ClassEval
ClassEval (Du et al., 2023) differs from the aforementioned datasets in terms of code generation scopes as it requires language models to generate class-level code passages rather than just a single function. It consists of 100 class-level code generation test examples in which a model has to generate a whole Python class code given a skeleton code of the class. Following the ClassEval paper, we use StarCoder-Instruct (GeorgiaTechResearchInstitute, 2023) as the base model and the same instruction-following template of the prompt used in the paper. As the entropy distribution from StarCoder-Instruct skews to lower values than that of the StarCoder model, we search hyperparameters for threshold in [0.01,0.03,0.05,0.1,0.2] and $\delta$ in [2,3,4,5,10,15,20]. We generate $n=5$ outputs to calculate pass@5. For EXP-edit, due to the high computational cost, we make 50 perturbations instead of 100.
E.3 Results
As presented in Table 2, SWEET outperforms all baselines in detecting machine-generated code in C++ and Java environments. Also, it still preserves code functionality much more than WLLM while achieving better detection performance. We observe that the Pareto Frontier lines of SWEET are ahead of those of WLLM, as in the Python environment (see Figure 7). It is worth noting that the C++ and Java examples in HumenEvalPack comprise longer code than the Python examples we used in the paper: an average of 100 tokens for C++, 97 tokens for Java, and 57 tokens for Python. Therefore, these results demonstrate that the efficacy of our methodology is not limited to the type of programming languages or the length of the code.
In the class-level code generation task, we could still observe our approach showing the highest AUROC score than the baselines even in the class-level code generation task where the LLM should generate longer and more complex code (the average token length of ClassEval solutions is 352). Specifically, SWEET achieves an AUROC of 0.980 with a 7.1% degradation of code functionality. Interestingly, as the text becomes longer, post-hoc methods’ performance increases. On the other hand, EXP-edit has shown lower detection performance than when evaluated in StarCoder, even in the high-entropy setting, due to the extremely spiky entropy distribution of StarCoder-Instruct. In addition, EXP-edit increases code functionality compared to non-watermarked baselines, even though it is a watermarking method that does not distort the original token distribution. We suppose these results are attributed to the specific watermark key sequence, which is randomly generated.